Hudaverdi Ltd pay current dividends of $0.96 per share with these dividends expected to grow at a rate of 4.6% per year in perpetuity. Hudaverdi Ltd shares are currently trading at $7.47 per share. What is the cost of equity finance for Hudaverdi Ltd? Give your answer as a percentage per annum to 1 decimal place.

Cost of equity =  % pa
Answer #1

Answer : Cost of equity (ke) = 18.04%

Calculation :

Ke = (D1 / P ) + g

Where,

D1 = Expected dividend ( i.e. 0.96 + 4.6% = 1.00416 )

P = Current price of share (i.e. 7.47)

G = growth rate (i.e. 4.6% or 0.046)

Ke = ( 1.00416 / 7.47 ) + 0.046

= 0.134425702811 + 0.046

= 0.180425702811 or 18.04%